OpenStudy (keigh2015):

If the area of a rectangle is 24a2b and the length is 8ab2, what would be the width of the rectangle, given that width is found by dividing area by length? Simplify the answer. the quantity of 3 times a all over b b over the quantity 3 times a 3 over the quantity a times b 3ab

OpenStudy (keigh2015):

@annas @Falling_In_Katt @uybuyvf @Poppies

OpenStudy (falling_in_katt):

3a/b

OpenStudy (anonymous):

divide like the problem says \[\frac{ 24a^2b }{ 8ab^2 }=\frac{ 24 }{ 8} \times \frac{ a^2 }{ a } \times \frac{ b }{ b^2 }\]

OpenStudy (keigh2015):

@undeadknight26

undeadknight26 (undeadknight26):

Divide: 24(a^2)b / 8a(b^2) = 24/8 * a^2/a * b/b^2 = ? Taken from: http://openstudy.com/study#/updates/52a26555e4b0a13bbbe93f8d